### v2

@@ 4.1.1 Collision probability
Figure 6: (a) Illustration of a rear-end collision (b) Rear-end collision probability due to robotic uncertainty.
@@ 4.2 Traffic states and the associated capacities
Figure 7: Traffic states consist of AVs’ desired trajectories under normal driving and with collisions.
@@ 4.3.1 The semi-Markov process
Since the stopping and restoring shock-waves are parallel , each point along the roadway segment will experience both the normal and abnormal states, with consistent durations for each state across all points, regardless of collision position. Furthermore, the dynamic changes in traffic states can be described by a semi-Markov process, as illustrated in Figure 8. The normal state is associated with a sojourn time of $\tau$, while the abnormal state has a sojourn time of $T$. For each point, the transition probability from the normal state to the abnormal state equals the collision rate on the roadway segment, $P(v,\eta)$, while the probability of remaining in the normal state is $1-P(v,\eta)$. Meanwhile, the abnormal state will certainly transition back to the normal state after time $T$. In addition, the collision rate $P(v,\eta)$ is assumed to be extremely small, as collision among AVs in fully autonomous traffic are rare events given the maturity of AV technology  [^3] However, theoretically, this value could exceed one, as it represents the expected number of vehicles involved in a collision on the roadway segment at each time step.. Therefore, the semi-Markov process governing traffic state changes remains mathematically well-defined.
@@
Figure 9: Relationship between macroscopic fully autonomous traffic mobility with desired speed and headway.
@@ 4.3.3 Comparison with collision-exclusive capacity
Figure 11: Comparison of our model with perfect operation assumption under different speeds and headway.
@@ 5.2.1 Capacity improvement
For any given maximum allowable collision probability $\hat{p}$ and fixed speed $v$, there exists a unique $\hat{\eta}_{v}$ such that
$p_{v}(\eta)\leq\hat{p}\iff\eta\geq\hat{\eta}_{v}.$ (14)
Furthermore, $\hat{\eta}_{v}$ satisfies $p_{v}(\hat{\eta}_{v})=\hat{p}$.
@@
Notice that $F_{d(v,\eta)}(l)=Prob(v\eta+\epsilon_{x}\leq l)$. This allows us to define $H_{v}(\eta;l)$ as a function of $\eta$, given by
$H_{v}(\eta;l)=Prob(\epsilon_{x}\leq l-v\eta)$
Clearly, $H_{v}(\eta;l)$ is a monotonically decreasing function of $\eta$. Therefore, $\eta$ and $p_{v}(\eta)$ a bijective mapping where $p_{v}(\eta)$ decreases with the increase of $\eta$
@@
When AVs’ operational speed is fixed at $v$, the original problem (13) can be formulated as
$\min_{\eta}\quad$ $\frac{1}{s_{v}(\eta)}=\eta+\frac{TL}{v\tau}p_{v}(\eta)$ (15)
$s.t.\quad$ $\eta\geq\hat{\eta}_{v}$
@@
The first order condition of the reformulated problem suggests that
$1+\frac{TL}{v\tau}p^{\prime}_{v}(\eta^{*})=0,$ (16)
where $\eta^{*}_{v}$ is the optimal solution of the underlying unconstrained program, and $p^{\prime}_{v}(\eta^{*}_{v})$ denotes the first order derivative, that is, $\frac{dp_{v}(\eta)}{d\eta}|_{\eta=\eta^{*}_{v}}$. As $p_{v}(\eta)$ is a monotonically decreasing function of $\eta$, it is theoretically possible that a $\eta^{*}_{v}\geq 0$ exists such that $p^{\prime}_{v}(\eta^{*}_{v})=-\frac{v\tau}{TL}$. Then, if $\forall\eta\geq\hat{\eta}_{v}$, $-\frac{v\tau}{TL}<p^{\prime}_{v}(\eta)<0$, the objective is monotonically increasing with $\eta$, making the optimal headway being $\hat{\eta}_{v}$. If $\eta^{*}_{v}\geq\hat{\eta}_{v}$ and $p^{\prime}_{v}(\eta^{*}_{v})=-\frac{v\tau}{TL}$, $\eta^{*}_{v}$ is the optimal headway. In conclusion, the optimal headway of problem (15) can be denoted as
$\eta^{{\dagger}}(v)=\max\{\hat{\eta}_{v},\eta^{*}_{v}\}.$ (17)
@@
Figure 13: The change of locally optimal headway according to different safety constraints.
@@
From Lemma 1, we know that $p(v,\eta)$ is monotonically decreasing with $\eta$. Using the same logic, we denote
$G_{\eta}(v;l)=Prob(\epsilon_{x}\leq l-v\eta)=H_{v}(\eta;l)$
which is a monotonically decreasing function of $v$, that is, $\frac{\partial p}{\partial v}<0$ Notice that
$dp(v,\eta)=\frac{\partial p}{\partial v}dv+\frac{\partial p}{\partial\eta}d\eta=0\Rightarrow\frac{d\eta}{dv}=-\frac{\partial p}{\partial v}/\frac{\partial p}{\partial\eta}$
@@
$=\frac{1}{\eta_{r}+\frac{TL}{\tau v_{2}}p(v_{2},\eta_{r})}$
$=\frac{1}{\eta_{r}+\frac{TL}{\tau v_{2}}p(v_{1},\eta^{{\dagger}}(v_{1}))}$
@@
$=s^{*}(v_{1})$
Here, the first inequality holds due to the optimality of $s_{\eta^{\dagger}}(v_{2})$. The following equality is derived by definition, while the second equality is based on the assumption in Eq (18). Then the last inequality stands since $\eta^{{\dagger}}(v_{1})>\eta_{r}$. The inequality implies that CIC under speed $v_{2}$ with optimal headway $\eta^{{\dagger}}(v_{2})$ is greater than that under speed $v_{1}$ with optimal headway $\eta^{{\dagger}}(v_{1})$. Since $v_{2}$ and $v_{1}$ are an $v_{2}>v_{1}$, the proof is concluded. ∎
@@
Figure 14: The general view of the optimization enlarging the capacity with strict and loose safety constraints.
@@ 5.2.2 Safety enhancement
We now claim that the CIC cannot increase infinitely with respect to $\eta$, which can be easily shown by contradiction. If not, as $\eta$ approaches infinity, we have
$\lim_{\eta\rightarrow+\infty}s=\frac{1}{\eta+\frac{TL}{\tau v}p(v,\eta)}=\frac{1}{+\infty+\frac{TL}{\tau v}p(v,\eta)}=+\infty$
This implies
$\lim_{\eta\rightarrow\infty}\frac{1}{s}=+\infty+\frac{TL}{\tau v}p(v,\eta)=0,$
which further suggests
$\lim_{\eta\rightarrow+\infty}\frac{TL}{\tau v}p(v,\eta)=-\infty.$
It then leads to the conclusion that
$\lim_{\eta\rightarrow+\infty}p(v,\eta)=-\infty\times\frac{\tau v}{TL}=-\infty.$
@@
For any pair of $v_{1}<v_{2}$, suppose the the corresponding optimal headway for problem (21) are $\eta^{r}_{v_{1}}$ and $\eta^{r}_{v_{2}}$. Since $s(v_{1},\eta^{r}_{v_{1}})=s(v_{2},\eta^{r}_{v_{2}})=\hat{s}$, we have
$\frac{1}{s(v_{2},\eta^{r}_{v_{1}})}=\eta^{r}_{v_{1}}+\frac{TL}{\tau v_{2}}p(v_{2},\eta^{r}_{v_{1}})$ $<\eta^{r}_{v_{1}}+\frac{TL}{\tau v_{1}}p(v_{1},\eta^{r}_{v_{1}})=\frac{1}{\hat{s}}=\eta^{r}_{v_{2}}+\frac{TL}{\tau v_{2}}p(v_{2},\eta^{r}_{v_{2}})$
The first inequality holds since the collision probability monotonically decreases with the increase of speed at the same headway. Given that Therefore, we derive that
$s(v_{2},\eta^{r}_{v_{1}})>s(v_{2},\eta^{r}_{v_{2}}).$ (22)
According to Lemma 4, $\eta^{r}_{v_{2}}$ is the highest $\eta$ that makes $s(v_{2},\eta^{r}_{v_{2}})\geq\hat{s}$. Therefore, $\forall\eta>\eta^{r}_{v_{2}}$, $s(v_{2},\eta)<\hat{s}$. In this context, we conclude from Eq (22) that
$\eta^{r}_{v_{1}}<\eta^{r}_{v_{2}}.$
Meanwhile, since $v_{1}<v_{2}$ by assumption, we further have
$p(v_{2},\eta^{r}_{v_{2}})<p(v_{2},\eta^{r}_{v_{1}})<p(v_{1},\eta^{r}{v_{1}}).$
@@
Figure 15: The general view of the optimization showing the safety enhancement with a demand constraint.
@@ 6.1 Overlapping collisions
In conclusion, the location of the second collision before the first one is cleared does not affect the nature of the capacity dynamics at each single location along the roadway, whose stationary property is captured by CIC introduced in the last section. Nonetheless, when considering the traffic throughput within a limited amount of time, the location of the second collision matters. In other words, the impacts of overlapping collisions and independent collisions are different. When a limited study period is assumed, we examine the expected value of traffic throughput of the roadway segment as the corresponding measure. Here, we provide its general form under overlapping collisions, $s_{c}(v,\eta)$ as follows:
@@
Here, $p_{c,i}$ shows the probability that the $i$th collision occurs, while $s_{c,i}$ indicates the associated throughput. In this context, $p_{c,0}$ represents the probability that no collision happens in the study period, and $s_{c,0}=s^{+}$. When the study period is assumed to be one hour, i.e., $H=1$ hr $=3600$, the specific form with two overlapping collisions is given by:
@@
$\approx\left(1-\frac{H}{\tau}\frac{L}{\eta v}p(v,\eta)+\frac{\frac{H}{\tau}\frac{L}{\eta v}\left(\frac{H}{\tau}\frac{L}{\eta v}-1\right)}{2}p(v,\eta)^{2}\right)s^{+}(\eta)$ (23c)
@@
$\approx\left(1-\left(\frac{H}{\tau}\frac{L}{\eta v}-1\right)p(v,\eta)\right)\frac{H}{\tau}\frac{L}{\eta v}p(v,\eta)\left(1-\frac{T}{H}\right)s^{+}(v,\eta)$ (23d)
Here, the equality in both equations holds due to the independence nature of the occurrence of collisions at different locations per time step. The approximation in Eq (23c) is based on binomial expansion since $p(v,\eta)$ is in general small, which condition is also used in the first approximation in Eq (23d). The second approximation in Eq (23d) is generated given that $p(v,\eta)^{3}$ is negligible.
The third term in Eq (23b) encloses two situations. In the first situation, the second collision happens in "Normal 1" and "Normal 2" of the first collision. We denote the corresponding probability and the remaining capacity as $p_{co,2}$ and $s_{co,2}$, respectively. The second situation depicts two independent collisions, whose corresponding probability and the remaining capacity as $p_{ci,2}$ and $s_{ci,2}$, respectively. As TCT is assumed to be more than half an hour, in this one hour example, two independent collisions will block the roadway, making the corresponding capacity $s_{ci,2}$ equal to 00 veh/hr. Together, the third term in Eq (23b) is given by:
@@
$=\frac{L(v+c)}{12vcH}\left(\frac{H}{\tau}\frac{L}{\eta v}\right)\left(\frac{H}{\tau}\frac{L}{\eta v}-1\right)p(v,\eta)^{2}\left(1-\frac{T}{H}\right)s^{+}(v,\eta)$ (23e)
@@
$s_{c}(v,\eta)\approx\left(1-\frac{T}{\tau}\frac{L}{\eta v}p(v,\eta)+\left(\frac{T}{H}-\frac{1}{2}+\frac{L(v+c)}{12vcH}\left(1-\frac{T}{H}\right)\right)\left(\frac{H}{\tau}\frac{L}{\eta v}\right)\left(\frac{H}{\tau}\frac{L}{\eta v}-1\right)p(v,\eta)^{2}\right)s^{+}(\eta)$ (23f)

### v3

@@ 4.1.1 Collision probability
Figure 6: (a) Illustration of a rear-end collision and the corresponding collision probability under stochastic car-following distance; (b) Rear-end collision probability due to robotic uncertainty with varying desired speed and headway.
@@ 4.2 Traffic states and the associated capacities
Figure 7: Traffic states consist of AVs’ desired trajectories under normal driving and with collisions (traffic trajectories are oriented from the bottom-left to the top-right).
@@ 4.3.1 The semi-Markov process
Since the stopping and restoring shock-waves are parallel, each point along the roadway segment will experience both the normal and abnormal states, with consistent durations for each state across all points, regardless of collision position. Furthermore, the dynamic changes in traffic states can be described by a semi-Markov process, as illustrated in Figure 8. The normal state is associated with a sojourn time of $\tau$, while the abnormal state has a sojourn time of $T$. For each point, the transition probability from the normal state to the abnormal state equals the collision rate on the roadway segment, $P(v,\eta)$, while that of remaining in the normal state is $1-P(v,\eta)$. Meanwhile, the abnormal state will certainly transition back to the normal state after time $T$. In addition, the collision rate $P(v,\eta)$ is assumed to be extremely small, as collision among AVs in fully autonomous traffic are rare events given the maturity of AV technology [^3] However, theoretically, this value could exceed one, as it represents the expected number of vehicles involved in a collision on the roadway segment at each time step.. Therefore, the semi-Markov process governing traffic state changes remains mathematically well-defined.
@@ 4.3.2 Discussions
Figure 9: Relationship between macroscopic fully autonomous traffic mobility with desired speed and headway. (a) Three-dimensional representation of CIC as a function of speed and headway; (b) Two-dimensional projection of CIC onto the speed-headway plane.
@@ 4.3.3 Comparison with collision-exclusive capacity
Figure 11: Comparison of our model results (blue curves) with perfect operation assumption (black curve) under different speeds and headway.
@@ 5.2.1 Capacity improvement
For any given maximum allowable collision probability $\hat{p}$ and fixed speed $v$, there exists a unique $\hat{\eta}_{v}$ such that $p_{v}(\eta)\leq\hat{p}\iff\eta\geq\hat{\eta}_{v}$. Furthermore, $\hat{\eta}_{v}$ satisfies $p_{v}(\hat{\eta}_{v})=\hat{p}$.
@@
Notice that $F_{d(v,\eta)}(l)=\textit{Prob}(v\eta+\epsilon_{x}\leq l)$. This allows us to define $H_{v}(\eta;l)$ as a function of $\eta$, given by:
$H_{v}(\eta;l)=\textit{Prob}(\epsilon_{x}\leq l-v\eta).$ (14)
Clearly, $H_{v}(\eta;l)$ is a monotonically decreasing function of $\eta$. Therefore, $\eta$ and $p_{v}(\eta)$ a bijective mapping where $p_{v}(\eta)$ decreases with the increase of $\eta$. ∎
@@
When AVs’ operational speed is fixed at $v$, the original problem (13) can be formulated as:
$\min_{\eta}\quad$ $\frac{1}{s_{v}(\eta)}=\eta+\frac{TL}{v\tau}p_{v}(\eta),$ (15a)
$s.t.\quad$ $\eta\geq\hat{\eta}_{v},$
@@
The first order condition of the reformulated problem suggests that:
$1+\frac{TL}{v\tau}p^{\prime}_{v}(\eta^{*})=0,$ (15b)
where $\eta^{*}_{v}$ is the optimal solution of the underlying unconstrained program, and $p^{\prime}_{v}(\eta^{*}_{v})$ denotes the first order derivative, that is, $\frac{dp_{v}(\eta)}{d\eta}|_{\eta=\eta^{*}_{v}}$. As $p_{v}(\eta)$ is a monotonically decreasing function of $\eta$, it is theoretically possible that a $\eta^{*}_{v}\geq 0$ exists such that $p^{\prime}_{v}(\eta^{*}_{v})=-\frac{v\tau}{TL}$. Then, if $\forall\eta\geq\hat{\eta}_{v}$, $-\frac{v\tau}{TL}<p^{\prime}_{v}(\eta)<0$, the objective is monotonically increasing with $\eta$, making the optimal headway being $\hat{\eta}_{v}$. If $\eta^{*}_{v}\geq\hat{\eta}_{v}$ and $p^{\prime}_{v}(\eta^{*}_{v})=-\frac{v\tau}{TL}$, $\eta^{*}_{v}$ is the optimal headway. In conclusion, the optimal headway of problem (15a) can be denoted as:
$\eta^{{\dagger}}(v)=\max\{\hat{\eta}_{v},\eta^{*}_{v}\}.$ (15c)
@@
Figure 13: The change of locally optimal headway according to different safety constraints The blue curve represents the CIC. The red line depicts the collision probability on the logarithmic scale. The light gray star indicates the optimal headway for $\hat{p}=10^{-8}$, and the dark gray star marks the optimal headway for $\hat{p}=10^{-10}$.
@@
From Lemma 1, we know that $p(v,\eta)$ is monotonically decreasing with $\eta$. Using the same logic, we denote:
$G_{\eta}(v;l)=\textit{Prob}(\epsilon_{x}\leq l-v\eta)=H_{v}(\eta;l),$ (16a)
which is a monotonically decreasing function of $v$, that is, $\frac{\partial p}{\partial v}<0$. Notice that:
$dp(v,\eta)=\frac{\partial p}{\partial v}dv+\frac{\partial p}{\partial\eta}d\eta=0\Rightarrow\frac{d\eta}{dv}=-\frac{\partial p}{\partial v}/\frac{\partial p}{\partial\eta}.$ (16b)
@@
$=\frac{1}{\eta_{r}+\frac{TL}{\tau v_{2}}p(v_{2},\eta_{r})}=\frac{1}{\eta_{r}+\frac{TL}{\tau v_{2}}p(v_{1},\eta^{{\dagger}}(v_{1}))}$
@@
$=s^{*}(v_{1}).$ (17b)
Here, the first inequality holds due to the optimality of $s_{\eta^{\dagger}}(v_{2})$. The following equality is derived by definition, while the second equality is based on the assumption in Eq (17a). Then the last inequality stands since $\eta^{{\dagger}}(v_{1})>\eta_{r}$. The inequality implies that CIC under speed $v_{2}$ with optimal headway $\eta^{{\dagger}}(v_{2})$ is greater than that under speed $v_{1}$ with optimal headway $\eta^{{\dagger}}(v_{1})$. Since $v_{2}$ and $v_{1}$ are an $v_{2}>v_{1}$, the proof is concluded. ∎
@@
Figure 14: Impact of safety constraints on the optimal CIC under varying speed. Contours show the CIC across speed–headway combinations. Blue curves Lines represent safety constraints and corresponding optimal headway for $\hat{p}\leq 10^{-8}$ and $\hat{p}\leq 10^{-10}$ with varying speed. Stars mark the maximum CIC under each constraint.
@@ 5.2.2 Safety enhancement
We now claim that the CIC cannot increase infinitely with respect to $\eta$, which can be easily shown by contradiction. If not, as $\eta$ approaches infinity, we have:
$\lim_{\eta\rightarrow+\infty}s=\frac{1}{\eta+\frac{TL}{\tau v}p(v,\eta)}=\frac{1}{+\infty+\frac{TL}{\tau v}p(v,\eta)}=+\infty,$ (20a)
that implies:
$\lim_{\eta\rightarrow\infty}\frac{1}{s}=+\infty+\frac{TL}{\tau v}p(v,\eta)=0,$ (20b)
which further suggests:
$\lim_{\eta\rightarrow+\infty}\frac{TL}{\tau v}p(v,\eta)=-\infty.$ (20c)
It then leads to the conclusion that:
$\lim_{\eta\rightarrow+\infty}p(v,\eta)=-\infty\times\frac{\tau v}{TL}=-\infty.$ (20d)
@@
For any pair of $v_{1}<v_{2}$, suppose the the corresponding optimal headway for problem (19) are $\eta^{r}_{v_{1}}$ and $\eta^{r}_{v_{2}}$. Since $s(v_{1},\eta^{r}_{v_{1}})=s(v_{2},\eta^{r}_{v_{2}})=\hat{s}$, we have:
$\frac{1}{s(v_{2},\eta^{r}_{v_{1}})}=\eta^{r}_{v_{1}}+\frac{TL}{\tau v_{2}}p(v_{2},\eta^{r}_{v_{1}})$ $<\eta^{r}_{v_{1}}+\frac{TL}{\tau v_{1}}p(v_{1},\eta^{r}_{v_{1}})=\frac{1}{\hat{s}}=\eta^{r}_{v_{2}}+\frac{TL}{\tau v_{2}}p(v_{2},\eta^{r}_{v_{2}}).$ (21a)
The first inequality holds since the collision probability monotonically decreases with the increase of speed at the same headway. Given that Therefore, we derive that:
$s(v_{2},\eta^{r}_{v_{1}})>s(v_{2},\eta^{r}_{v_{2}}).$ (21b)
According to Lemma 4, $\eta^{r}_{v_{2}}$ is the highest $\eta$ that makes $s(v_{2},\eta^{r}_{v_{2}})\geq\hat{s}$. Therefore, $\forall\eta>\eta^{r}_{v_{2}}$, $s(v_{2},\eta)<\hat{s}$. In this context, we conclude from Eq (21b) that:
$\eta^{r}_{v_{1}}<\eta^{r}_{v_{2}}.$ (21c)
Meanwhile, since $v_{1}<v_{2}$ by assumption, we further have:
$p(v_{2},\eta^{r}_{v_{2}})<p(v_{2},\eta^{r}_{v_{1}})<p(v_{1},\eta^{r}{v_{1}}).$ (21d)
@@
Figure 15: Safety enhancement with a demand constraint. The blue line confines the feasible region under demand constraint. The star symbol marks the corresponding CIC when collision probability is minimized.
@@ 6.1 Overlapping collisions
In conclusion, the location of the second collision before the first one is cleared does not affect the nature of the capacity dynamics at each single location along the roadway, whose stationary property is captured by CIC introduced in the last section. Nonetheless, when considering the traffic throughput within a limited amount of time, the location of the second collision matters. In other words, the impacts of overlapping collisions and independent collisions are different. When a limited study period is assumed, we examine the expected value of traffic throughput of the roadway segment as the corresponding measure. Here, we provide the general form of CIC for one lane under overlapping collisions, $s_{c}(v,\eta)$, as follows:
@@
Here, $p_{c,i}$ shows the probability that the $i$th collision occurs, while $s_{c,i}$ indicates the associated throughput. In this context, $p_{c,0}$ represents the probability that no collision happens in the study period, and $s_{c,0}=s^{+}$. When the study period is assumed to be one hour, i.e., $H=1$ hr $=3600$ s, the specific form with at most two overlapping collisions is given by:
@@
$\approx\left(1-\frac{H}{\tau}\frac{L}{\eta v}p(v,\eta)+\frac{\frac{H}{\tau}\frac{L}{\eta v}\left(\frac{H}{\tau}\frac{L}{\eta v}-1\right)}{2}p(v,\eta)^{2}\right)s^{+}(\eta);$ (22c)
@@
$\approx\left(1-\left(\frac{H}{\tau}\frac{L}{\eta v}-1\right)p(v,\eta)\right)\frac{H}{\tau}\frac{L}{\eta v}p(v,\eta)\left(1-\frac{T}{H}\right)s^{+}(v,\eta).$ (22d)
Here, the equalities in both equations holds due to the independence nature of the occurrence of collisions at different locations within time step. Specifically, the probability that no collisions occur during the study period, denoted by $p_{c,0}(v,\eta)$, is given by to $\left(1-p(v,\eta)\right)^{\frac{H}{\tau}\frac{L}{\eta v}}$, and the associated capacity under this no-collision scenario corresponds to the full capacity $s^{+}(v,\eta)$. The approximation in Eq (22c) is based on a binomial expansion, which is justified by the fact that $p(v,\eta)$ is in general small. The second term accounts for the scenario in which exactly one collision occurs during the study. This occurs with a probability of $\frac{H}{\tau}\frac{L}{\eta v}p(v,\eta)\left(1-p(v,\eta)\right)^{\frac{H}{\tau}\frac{L}{\eta v}-1}$, and the corresponding throughput is reduced to $\left(1-\frac{T}{H}\right)s^{+}(v,\eta)$. A binomial expansion is also applied to the first approximation in Eq (22d), and the second approximation holds under the assumption that $p(v,\eta)$ is sufficiently small such that higher-order terms beyond $p(v,\eta)^{2}$ can be negligible.
The third term in Eq (22b), representing the two-accidents scenario, encloses two detailed situations. In the first situation, the two collisions happens in the "Normal 1" or "Normal 2" regions of the original collision (see Figures 16 (a) and (b)) and thereby are overlapping. We denote the corresponding probability and the remaining capacity as $p_{co,2}$ and $s_{co,2}$, respectively. As previously analyzed, its remaining capacity is equivalent to that under the one-collision scenario, that is, $s_{co,2}=\left(1-\frac{T}{H}\right)s^{+}(v,\eta)$.
The second situation depicts two independent collisions, whose corresponding probability and the remaining capacity are represented as $p_{ci,2}$ and $s_{ci,2}$, respectively. As TCT is assumed to exceed half an hour, two independent collisions within the study period, which is set to one hour, would completely block the roadway, making the corresponding capacity $s_{ci,2}$ equal to 00 veh/hr. Together, the third term in Eq (22b) is given by:
@@
$=\frac{L(v+c)}{12vcH}\left(\frac{H}{\tau}\frac{L}{\eta v}\right)\left(\frac{H}{\tau}\frac{L}{\eta v}-1\right)p(v,\eta)^{2}\left(1-\frac{T}{H}\right)s^{+}(v,\eta).$ (22e)
Similarly, the approximation is also based on binomial expansion with terms with equal or higher order than $p(v,\eta)^{2}$ since $p(v,\eta)$ is small.
@@
$s_{c}(v,\eta)\approx\left(1-\frac{T}{\tau}\frac{L}{\eta v}p(v,\eta)+\left(\frac{T}{H}-\frac{1}{2}+\frac{L(v+c)}{12vcH}\left(1-\frac{T}{H}\right)\right)\left(\frac{H}{\tau}\frac{L}{\eta v}\right)\left(\frac{H}{\tau}\frac{L}{\eta v}-1\right)p(v,\eta)^{2}\right)s^{+}(\eta).$ (22f)